# Experimental Demonstration of a Decentralized Electromagnetic Formation Flying Control Using Alternating Magnetic Field Forces

### v1

@@ 3 Simulation
$P$ $02\times 10^{-4} 2.2$$27\times 10^{-4}$ $1.43\times 10^{-4}$ $1.41\times 10^{-4}$
@@
This experiment demonstrates closed-loop attraction. The satellites start at a relative distance of approximately $0.5$ m. The objective is to reach a desired relative distance $|d_{ij}|=0.45$ m. Figure 8 shows that $|\hat{r}_{ij}|$ settles to the desired relative position $|d_{ij}|=0.45$ m after approximately 20 s, which demonstrates that the satellites achieve formation. Note that the satellites move before the control is turned on at $t=5$ s. This can be explained by a slight impulse introduced at the start of the experiment and airflow from the airtracks creating disturbances.
@@ 4 Simulation
$P$ $2.38\times 10^{-4}$ $2.62\times 10^{-4}$ $1.66\times 10^{-4}$ $1.64\times 10^{-4}$
@@
Experiment 5.
This experiment demonstrates multiple maneuvers in succession. The satellites start at a relative distance of approximately 0.50 m. The initial objective is to reach a desired relative distance $|d_{ij}|=0.4$ m. Next, the desired relative position is switched to $0.5$ m at $t=60$ s. Figure 10 shows that $|\hat{r}_{ij}|$ settles to $|d_{ij}|=0.40$ m at approximately 35 s, and $|\hat{r}_{ij}|$ settles to $|d_{ij}|=0.50$ m by approximately 90 s, which demonstrates that the satellites can achieve multiple formation maneuvers. Figure 11 shows the simulation results, which qualitatively matches the experimental results in Figure 10. $\triangle$
7.2 3 Satellite Closed-Loop Experiments
Next, we present 3 closed-loop experiments with $n=3$ satellites. The middle, left, and right satellites are numbered 1, 2, and 3. The neighbor sets are $\mathcal{N}_{1}=\{2,3\}$, $\mathcal{N}_{2}=\{1\}$, and $\mathcal{N}_{3}=\{1\}$. The interaction frequencies are $\omega_{12}=\omega_{21}=20\pi$ rad/s and $\omega_{13}=\omega_{31}=40\pi$ rad/s. We implement the currents $(I_{ij,k},I_{ji,k})$ given by (29)–(30) with $\epsilon_{0}=0.015$ m, $\epsilon_{1}=0.021$ m, $\alpha_{ij}=0.0158$, $\beta=7.38$, $\rho_{ij}=0.00136$, $\gamma_{12}=\gamma_{13}=0.8$ and $\gamma_{21}=\gamma_{31}=1.25$, and $\bar{I}=2.35$ A. Note that we select $\alpha_{ij}$, $\beta$, and $\rho_{ij}$ such that $\alpha_{ij}m/(c_{0}N^{2}A^{2})=812$, $\alpha_{ij}\beta m/(c_{0}N^{2}A^{2})=6000$, and $\rho_{ij}m/(c_{0}N^{2}A^{2})=70$.
Using experimental data, we estimate the noise variances $V=2\times 10^{-6}$ $\mathrm{m}^{2}$ and $w=5\times 10^{-6}$ $\mathrm{m}^{2}/\mathrm{s}^{4}$. It follows from 36, 37, and 38 that $L=\begin{bmatrix}0.1064&0.0598\end{bmatrix}^{\mathrm{T}}$.
@@
Experiment 6.
This experiment demonstrates closed-loop repulsion between satellite 1 and 2, and closed-loop repulsion between satellite 1 and 3. Satellites 1 and 2 start at a relative distance of approximately 0.35 m, and satellites 1 and 3 start at a relative distance of approximately 0.38 m apart. The objective is to reach the desired relative positions $d_{12}=0.42$ m and $d_{13}=-0.45$ m. Figure 12 shows the experimental results, where the column on the left shows relative measurements, estimates, and average approximate force between satellite 2 (left satellite) and satellite 1 (middle satellite), as measured and computed on satellite 2; and the second column from the left shows the data as measured and computed on satellite 1. The third column from the left shows relative measurements, estimates, and average approximate force between satellite 1 and satellite 3 (right satellite), as measured and computed on satellite 1; and the column on the right shows the data as measured and computed on satellite 3. Figure 12 shows that $\hat{r}_{12}$ and $\hat{r}_{13}$ converges to $d_{12}$ and $d_{13}$ at approximately $t=30$ s.
Next, we present simulations of the closed-loop 3 and 4 where $\mathbf{p}_{ij,k}=p_{ij,k}\mathbf{i}$. The effect of friction between the satellites and the linear air tracks becomes noticeable with the addition of the third satellite. To model this effect, we include linear damping $\hat{B}_{i}v_{i}$ in 3, where $\hat{B}_{i}=0.08$ N–s/m. The simulated relative position includes zero-mean Gaussian-white sensor noise with variance $V=2\times 10^{-6}$ m2, which is the sensor noise variance estimated from experimental data. Figure 13 shows the simulation results, which qualitatively agree with the experimental results. However, we observe an oscillatory transient response in Figure 13, which is not observed in Figure 12. This can partly be explained by the imperfect friction model, unmodeled external disturbances (e.g., lateral air pressure from the air track), and unmodeled near-field electromagnetic effects. $\triangle$
Experiment 7.
This experiment demonstrates closed-loop attraction between satellite 1 and satellite 2, and closed-loop attraction between satellite 1 and satellite 3. Satellites 1 and 2 start at a relative distance of approximately 0.425 m, and satellites 1 and 3 start at a relative distance of approximately 0.46 m. The objective is to reach the desired relative positions $d_{12}=0.35$m and $d_{13}=-0.38$ m. Figure 14 shows that $\hat{r}_{12}$ and $\hat{r}_{13}$ converges to the desired relative positions $d_{12}$ and $d_{13}$ by approximately $t=26$ s, which demonstrates that the satellites achieve formation. Figure 15 shows the simulation results, which qualitatively agree with the experimental results. Similar to the comparison in 6, the simulation has oscillatory overshoot whereas the overshoot in the experiment decays more quickly. $\triangle$
@@
This experiment demonstrates closed-loop attraction between satellite 1 and satellite 2, and closed-loop repulsion between satellite 1 and satellite 3. Satellites 1 and 2 start at a relative distance of approximately 0.46 m, and satellites 1 and 3 start at a relative distance of approximately 0.38 m apart.The objective is to reach the desired relative positions $d_{12}=0.42$ m and $d_{13}=-0.45$ m. Figure 16 shows that $\hat{r}_{12}$ and $\hat{r}_{13}$ converges to the desired relative position $d_{12}$ and $d_{13}$ by approximately $t=25$ s. Figure 17 shows the simulation of the experiment, which qualitatively matches the experimental results in Figure 16. $\triangle$
@@ 8 Concluding Remarks
This article presented a 3-satellite experimental demonstration of decentralized EMFF using AMFF. To our knowledge, this is the first demonstration of AMFF with at least 3 satellites in open or closed loop. The experimental results in this article not only demonstrated the feasibility of EMFF with AMFF but also demonstrated that the frequency-multiplexed amplitude-modulated sinusoidal approach can be used to achieve formation with multiple satellites. Future experiments could include 3 degree-of-freedom experiments (planar translation and attitude control) for 2 or more satellites.
@@ Acknowledgments
Appendix A Proof of Proposition 2
First, consider (i) $\mathbf{r}\times\mathbf{f}_{*}\neq\mathbf{0}$, and substituting $|\mathbf{r}\times\mathbf{f}_{*}|=|\mathbf{r}|^{2}\mathbf{f}_{*}-(\mathbf{r}\cdot\mathbf{f}_{*})\mathbf{r}$, (16), and (17) into (2) we get
$g_{\mathrm{r}}h_{\mathrm{rf}}$ $=g_{\mathrm{rf}}h_{\mathrm{r}}=\frac{|\mathbf{r}\times\mathbf{f}_{*}|}{2|\mathbf{r}|}$
$a_{y}b_{y}$ $=-\frac{\operatorname{sgn}(\mathbf{r}\cdot\mathbf{f}_{*})}{2|\mathbf{r}|}\mathopen{}\mathclose{{\left(-|\mathbf{r}\cdot\mathbf{f}_{*}|+\Phi_{1}}}\right),$
Next, we consider (ii) $\mathbf{r}\times\mathbf{f}_{*}=\mathbf{0}$, and substituting (10)– (13), (16), and (17) into (2) yields

### v2

@@ 3 Simulation
$P_{\hat{F}}$ (N) $02\times 10^{-4} 2.2$$27\times 10^{-4}$ $1.43\times 10^{-4}$ $1.41\times 10^{-4}$
@@
This experiment demonstrates closed-loop attraction. The satellites start at a relative distance of approximately $0.5$ m. The objective is to reach a desired relative distance $|d_{ij}|=0.45$ m. Figure 9 shows that $|\hat{r}_{ij}|$ settles to $|d_{ij}|=0.45$ m after approximately 20 s. Specifically, the settling time is $21.0$ s, and the mean steady-state error $\hat{r}_{ij}-d_{ij}$ over the last 60 s is $3.8\times 10^{-3}$ m. Note that the satellites move before the control is turned on at $t=5$ s. This can be explained by a slight impulse introduced at the start of the experiment and airflow from the airtracks creating disturbances. The variance of the unfiltered steady-state error $q_{ij}-d_{ij}$ is $3.3\times 10^{-6}$ m2, which is comparable to 3.
@@ 4 Simulation
$P_{\hat{F}}$ (N) $2.38\times 10^{-4}$ $2.62\times 10^{-4}$ $1.66\times 10^{-4}$ $1.64\times 10^{-4}$
Experiment 5.
This experiment demonstrates multiple maneuvers in succession. The satellites start at a relative distance of approximately 0.50 m. The initial objective is to reach a desired relative distance $|d_{ij}|=0.4$ m. Next, the desired relative position is switched to $0.5$ m at $t=60$ s. Figure 11 shows that $|\hat{r}_{ij}|$ settles to $|d_{ij}|=0.40$ m at approximately 35 s, and $|\hat{r}_{ij}|$ settles to $|d_{ij}|=0.50$ m by approximately 90 s, which demonstrates that the satellites can achieve multiple formation maneuvers. The current amplitude saturation 32 is active from $5.1$ s to $6.3$ s. During this time interval, the unsaturated current amplitudes exceed the limit $\bar{I}=2.35$ A. The peak unsaturated amplitude is $2.77$ A, which occurs at $5.3$ s and corresponds to a desired force of $0.0078$ N. Since the amplitude is saturated, the time-averaged force is only $0.0062$ N, which is 20% less than the desired value. This demonstrates that the control approach has some robustness to the current amplitude limits imposed by the hardware. Figure 12 shows the simulation results, which qualitatively matches the experimental results in Figure 11. $\triangle$
9.2 Three-Satellite Closed-Loop Formation Experiments
Next, we present 3 closed-loop experiments with $n=3$ satellites. The middle, left, and right satellites are numbered 1, 2, and 3. The neighbor sets are $\mathcal{N}_{1}=\{2,3\}$, $\mathcal{N}_{2}=\{1\}$, and $\mathcal{N}_{3}=\{1\}$. The interaction frequencies are selected as the 2 smallest common multiples of $2\pi/T$, where $T=0.1$ s. Specifically, $\omega_{12}=\omega_{21}=20\pi$ rad/s and $\omega_{13}=\omega_{31}=40\pi$ rad/s. The smallest common multiples are used because lower frequency results in smaller impedance and larger maximum allowable current. We implement the currents $(I_{ij,k},I_{ji,k})$ given by (31)–(32) with $\epsilon_{0}=0.015$ m, $\epsilon_{1}=0.021$ m, $\alpha_{ij}=0.0158$, $\beta=7.38$, $\rho_{ij}=0.00136$, $\gamma_{12}=\gamma_{13}=0.8$ and $\gamma_{21}=\gamma_{31}=1.25$, and $\bar{I}=2.35$ A. Note that we select $\alpha_{ij}$, $\beta$, and $\rho_{ij}$ such that $\alpha_{ij}m/(c_{0}N^{2}A^{2})=812$, $\alpha_{ij}\beta m/(c_{0}N^{2}A^{2})=6000$, and $\rho_{ij}m/(c_{0}N^{2}A^{2})=70$. The gains $\alpha_{ij}$, $\beta$, and $\rho_{ij}$ are selected based on simulation responses. Using experimental data, we estimate the variances $V=2\times 10^{-6}$ $\mathrm{m}^{2}$ and $w=5\times 10^{-6}$ $\mathrm{m}^{2}/\mathrm{s}^{4}$. It follows from 38, 39, and 40 that
$P=\begin{bmatrix}0.3891&0.3456\\ 0.3456&0.5879\end{bmatrix}\times 10^{-6},\qquad L=\begin{bmatrix}0.1773\\ 0.1447\end{bmatrix}$.
@@ 9.2 Three-Satellite Closed-Loop Formation Experiments
Experiment 6.
This experiment demonstrates closed-loop repulsion between satellites 1 and 2, and closed-loop repulsion between satellites 1 and 3. Satellites 1 and 2 start at a relative distance of approximately 0.35 m, and satellites 1 and 3 start at a relative distance of approximately 0.38 m apart. The objective is to reach the desired relative positions $d_{12}=0.42$ m and $d_{13}=-0.45$ m. Figure 13 shows the experimental results, where the column on the left shows relative measurements, estimates, and time-averaged approximate force between satellite 2 (left satellite) and satellite 1 (middle satellite), as measured and computed on satellite 2; and the second column from the left shows the data as measured and computed on satellite 1. The third column from the left shows relative measurements, estimates, and time-averaged approximate force between satellite 1 and satellite 3 (right satellite), as measured and computed on satellite 1; and the column on the right shows the data as measured and computed on satellite 3. Figure 13 shows that $\hat{r}_{12}$ and $\hat{r}_{13}$ converges to $d_{12}$ and $d_{13}$ after approximately $26$ s. For satellites 1 and 2, the settling time is 28.3 s, and the mean steady-state error $\hat{r}_{12}-d_{12}$ over the last 60 s is $1.0\times 10^{-3}$ m. For satellites 1 and 3, the settling time is 26.4 s, and the mean steady-state error $\hat{r}_{13}-d_{13}$ over the last 60 s is $0.9\times 10^{-3}$ m. The variance of the unfiltered steady-state error $q_{ij}-d_{ij}$ is $4.1\times 10^{-6}$ m2, where $2\times 10^{-6}$ is attributed to sensor noise variance and the remaining is potentially due in part to unmodeled external disturbances.
Next, we present simulations of the closed-loop 3 and 4 where $\mathbf{p}_{ij,k}=p_{ij,k}\mathbf{i}$. The effect of friction between the satellites and the linear air tracks becomes noticeable with the addition of the third satellite. This effect is modeled by including linear damping $b_{i}v_{i}\mathbf{i}$ in 3, where selecting $b_{i}$ involves a tradeoff between reducing response oscillations and increasing settling time. Larger $b_{i}$ tends to decrease oscillations but increase settling time. The damping coefficient $b_{i}=0.08$ N-s/m is selected to achieve a similar response to the experiments. The simulated relative position includes zero-mean Gaussian-white sensor noise with variance $V=2\times 10^{-6}$ m2, which is the sensor noise variance estimated from experimental data. Figure 14 shows the simulation results, which qualitatively agree with the experimental results. However, we observe an oscillatory transient response in Figure 14, which is not observed in Figure 13. This can partly be explained by the imperfect friction model, unmodeled external disturbances (e.g., lateral air pressure from the air track), and unmodeled near-field electromagnetic effects. $\triangle$
@@
Experiment 7.
This experiment demonstrates closed-loop attraction between satellite 1 and satellite 2, and closed-loop attraction between satellite 1 and satellite 3. Satellites 1 and 2 start at a relative distance of approximately 0.425 m, and satellites 1 and 3 start at a relative distance of approximately 0.46 m. The objective is to reach the desired relative positions $d_{12}=0.35$m and $d_{13}=-0.38$ m. Figure 15 shows that $\hat{r}_{12}$ and $\hat{r}_{13}$ converges to the desired relative positions $d_{12}$ and $d_{13}$ by approximately $26$ s. For satellites 1 and 2, the settling time is 23.1 s, and the mean steady-state error $\hat{r}_{12}-d_{12}$ over the last 60 s is $0.7\times 10^{-3}$ m. For satellites 1 and 3, the settling time is 25.5 s, and the mean steady-state error $\hat{r}_{13}-d_{13}$ over the last 60 s is $0.8\times 10^{-3}$ m. The variance of the unfiltered steady-state error $q_{ij}-d_{ij}$ is $4.1\times 10^{-6}$ m2, which is comparable to 6. Figure 16 shows the simulation results, which qualitatively agree with the experimental results. Similar to the comparison in 6, the simulation has oscillatory overshoot whereas the overshoot in the experiment decays more quickly. $\triangle$
@@
This experiment demonstrates closed-loop attraction between satellite 1 and satellite 2, and closed-loop repulsion between satellite 1 and satellite 3. Satellites 1 and 2 start at a relative distance of approximately 0.46 m, and satellites 1 and 3 start at a relative distance of approximately 0.38 m apart. The objective is to reach the desired relative positions $d_{12}=0.42$ m and $d_{13}=-0.45$ m. Figure 17 shows that $\hat{r}_{12}$ and $\hat{r}_{13}$ converges to the desired relative position $d_{12}$ and $d_{13}$ by approximately $25$ s. For satellites 1 and 2, the settling time is 24.6 s, and the mean steady-state error $\hat{r}_{12}-d_{12}$ over the last 60 s is $1.2\times 10^{-3}$ m. For satellites 1 and 3, the settling time is 24.4 s, and the mean steady-state error $\hat{r}_{13}-d_{13}$ over the last 60 s is $1.6\times 10^{-3}$ m. The variance of the unfiltered steady-state error $q_{ij}-d_{ij}$ is $4.0\times 10^{-6}$ m2, which is comparable to 6 and 7. Figure 18 shows the simulation of the experiment, which qualitatively matches the experimental results in Figure 17. $\triangle$
@@
Table 3 is a summary of experimental results, including the settling time, mean steady-state error $\hat{r}_{ij}-d_{ij}$ over the last 60 s, variance of the steady-state error over last 60 s, and the maximum steady-state error over last 60 s.
Table 3: Settling time, mean steady-state error, variance of steady-state error, and maximum steady-state error
Mean steady- Maximum steady-
Experiment $T_{{\rm s}}$ (s) state error (m) Variance (m2) state error (m)
3 21.1 $2.7\times 10^{-3}$ 8.1 $\times 10^{-7}$ $6.0\times 10^{-3}$
4 21.0 $3.8\times 10^{-3}$ 13.7 $\times 10^{-7}$ $8.3\times 10^{-3}$
6 27.3 $0.95\times 10^{-3}$ 7.4 $\times 10^{-7}$ $4.1\times 10^{-3}$
7 24.3 $0.75\times 10^{-3}$ 3.9 $\times 10^{-7}$ $4.0\times 10^{-3}$
8 24.5 $1.4\times 10^{-3}$ $5.2\times 10^{-7}$ $5.4\times 10^{-3}$
@@ 10 Concluding Remarks
This article presented a 3-satellite experimental demonstration of decentralized EMFF using AMFF. To the authors’ knowledge, this is the first demonstration of AMFF with at least 3 satellites in open or closed loop. The experimental results not only demonstrate the feasibility of EMFF with AMFF but also demonstrate that the frequency-multiplexed amplitude-modulated sinusoidal approach can be used to achieve formation with multiple satellites. Specifically, the closed-loop experiments demonstrate decentralized EMFF with AMFF for 3 satellites where the mean steady-state formation error is less than $0.005$ m, the maximum steady-state formation error is less than $\pm 0.01$ m, and the settling time is less than $30$ s. The experiments also show that the approach in this paper provides robustness to hardware limitations (e.g., power limits, sensor noise). However, future experimental studies could adopt the AMFF control methods in [10] to explicitly account for constraints (e.g., power/current limits, thermal management of the coils) using a composite control barrier function.
Future experimental studies could focus on considerations that are relevant for in-space testing and deployment. Although the method in this article applies to 3-degree-of-freedom position control, the experiments are restricted to a single degree of freedom. Future experiments could include ground-based 3-degree-of-freedom experiments (planar translation and one-degree-of-freedom attitude control). Such experiments would require multiple electromagnetic coils (e.g., 2 orthogonal coils for control in the plane) and additional interaction frequencies for relative attitude control [2]. Similarly, 6-degree-of-freedom would require at least 3 electromagnetic coils. In this case, hardware design is tightly interlinked with power requirements, computational complexity of the control algorithms, and acceptable performance levels.
Other important considerations for on-orbit deployment include accounting for gravitational forces (which is addressed in [10]), disturbances such as atmospheric drag in low Earth orbit, and other space environment effects (e.g., radiation, extreme heat-cold cycles, Earth’s magnetic field).